OpenStudy (anonymous):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits.

OpenStudy (isaiah.feynman):

If \[l = 6.25\] then what is \[l^{2}\]?

OpenStudy (anonymous):

(6 + 1/4)^2 = 6^2 + 2*6*(1/4) + (1/4)^2 = 36 + 3 + 1/16 = 36 + 3 + .0625

OpenStudy (isaiah.feynman):

@pgpilot326 I like your method with fractions!

OpenStudy (anonymous):

thanks @Isaiah.Feynman !